OpenStudy (anonymous):

solve using substitution 7X-3Y=20 5X+3Y=16

OpenStudy (lucaz):

isolate one of the variables in one equation and substitute on the other equation

OpenStudy (lucaz):

before you do that, you can add the equations and get 12x+0=36

OpenStudy (lucaz):

the x = 36/12 = 3 then you can substitute it on the equation

OpenStudy (anonymous):

Thanks I already figured it out though...l